# Uniformly perfect measures on strictly convex planar graphs are $L^{2}$-flattening

### v1

@@ Abstract.
Fourier decay, inverse theorems, sumsets, convolutions, energy.
@@ 1. Introduction
The following corollary explains why Theorem 1.2 is an $L^{2}$-flattening result:
Corollary 1.3.
For every $D>1$, $\beta\in[0,1)$, $\mathfrak{d}>0$, and $\epsilon\in(0,1)$ there exists $\kappa=\kappa(D,\beta,\epsilon)>0$ such that the following holds.
Let $R\geq 1$, $f\in L^{2}(\mathbb{R}^{2})$, and $\sigma$ a Borel probability measure such that:
  1. (1)
    $\operatorname{spt}\hat{f}\subset B(R)$ and $\|f\|_{L^{2}}\geq R^{\epsilon}\|f\|_{L^{1}}$.
  2. (2)
    $\sigma$ is a $(D,\beta)$-uniformly perfect measure on $\mathbb{P}$ such that $\operatorname{diam}(\operatorname{spt}\sigma)\geq\mathfrak{d}$.
Then,
$\|f\ast\sigma\|_{L^{2}}\lesssim_{D,\mathfrak{d},\beta,\eta}R^{-\kappa}\|f\|_{L^{2}}.$
The derivation of Corollary 1.3 from Theorem 1.2 is similar to the proof of [2, Corollary 1.2], where the same phenomenon is phrased in slightly different notation. We nonetheless provide the details in Section 4.3.
@@ 1.1. Previous work
Close relatives of Theorem 1.2 in previous literature are the $L^{2}$-flattening theorems of Rossi-Shmerkin [25, Theorem 1.1] on $\mathbb{R}$, and Khalil [17, Theorem 1.6] on $\mathbb{R}^{d}$. These results are formulated in terminology different from Theorem 1.2, but they are roughly the counterparts of Theorem 1.2 for (i) uniformly perfect measures on $\mathbb{R}$, and (ii) measures on $\mathbb{R}^{d}$ satisfying Khalil’s uniformly affine non-concentration condition. The reader should note that measures supported on smooth curves – as in Theorem 1.2 – do not satisfy Khalil’s non-concentration condition, due to the presence of tangent lines; otherwise Theorem 1.2 could be deduced from the results in [17].
Thus, Theorem 1.2 and [17, Theorem 1.6] are complementary, but finding a (natural) common generalisation seems like an interesting problem. For expert readers, we also mention that existing higher-dimensional inverse theorems (by Hochman [16] and Shmerkin [27]) do not appear to be powerful enough to prove Theorem 1.2, again due to the existence of tangent lines. Our proof will eventually rely on the one-dimensional inverse theorem of Shmerkin [26, Theorem 2.1], see Section 1.2 for a brief explanation.
@@
Finally, Theorem 1.2 is related to several recent works studying Fourier decay of stationary measures. The first author and Khalil [2] proved that the conclusion of Theorem 1.2 holds under the following assumptions. The measure $\sigma$ is the lift of a non-atomic self-similar measure on $\mathbb{R}$, onto (a) either an analytic curve whose trace is not contained in an affine hyperplane of $\mathbb{R}^{d}$, or (b) a $C^{d+1}$-curve $\gamma$ such that $\{\gamma^{\prime},\gamma^{\prime\prime},...,\gamma^{(d)}\}$ span $\mathbb{R}^{d}$ at every point. This was the first demonstration of a general non-trivial class of measures on curves for which $L^{2}$ flattening can be obtained, in the sense (1.3). This work is also related to that of Algom, Chang, Meng Wu, and Yu-Liang Wu [1], and the simultaneous independent work by Baker and Banaji [6], on pointwise Fourier decay for smooth strictly convex push-forwards of self-similar measures. It is also related to the subsequent paper of Baker, Khalil, and Sahlsten [7], as well as to [3, 4, 8, 9, 11, 19, 28, 29, 30]. For more on the relation between Theorem 1.2 and the Fourier decay problem for stationary measures, see [2, Section 1.2].
@@
It is also natural to consider analogues of Theorem 1.2 for surfaces. For instance, one may ask whether a Borel probability measure $\sigma$ on the paraboloid (or any other "curved" hypergraph) in $\mathbb{R}^{d+1}$ is $L^{2}$ flattening, provided its projection to $\mathbb{R}^{d}$ is uniformly affinely non-concentrated in the sense of Khalil [17]. More generally, it may be plausible that $\sigma$ is $L^{2}$ flattening whenever its projection to $\mathbb{R}^{d}$ is $L^{2}$ flattening. A strategy like this underlies the proof of $L^{2}$ flattening for self-similar measures on curves, [2, Theorem 1.3].
@@ 1.2. Proof outline
Our proof is based on Shmerkin’s inverse theorem for $L^{q}$-norms [26, Theorem 2.1], which in turn is inspired by Hochman’s inverse theorem for entropy [15, Theorem 2.7]. In fact, we employ a corollary of Shmerkin’s theorem obtained by Rossi and Shmerkin [25, Proposition 3.1], which states that convolution with a uniformly perfect measure on the real line is $L^{2}$-flattening. One may think of $Y$ in the "Lemma" as the support of our uniformly perfect measure $\sigma$ in Theorem 1.2, although this is slightly misleading; Lemma 3.2 deals directly with the measure $\sigma$, not its support. The support analogy is more accurate if $\sigma$ happens to be Ahlfors regular (and Figure 1 depicts this case).
@@
From Proposition 3.3 (and a "measure-theoretic" version of it stated in Proposition 3.1) adapting a strategy developed in [21, 22], we derive Proposition 4.2: this statement could be summarised by saying that convolution with a uniformly perfect measure $\sigma$ on $\mathbb{P}$ is "Riesz energy flattening". In particular, taking repeated self-convolutions of $\sigma$ gradually increases the highest index $\alpha$" for which the $\alpha$-dimensional Riesz energy of $\sigma^{k}$ is (nearly) bounded. This is formalised in Corollary 4.6. Eventually, the conclusion is that the $(2-\epsilon)$-dimensional Riesz energy of $\sigma^{k}$ is (nearly) bounded, and this easily yields Theorem 1.2, see Section 4.2.
@@ 2.2. Rossi and Shmerkin’s theorem
We need this refined definition since in our application of Rossi and Shmerkin’s theorem, we shall require certain restrictions of the original measure to be uniformly perfect (see Lemma 2.14 below). Now, by [25, Lemma 4.1], Ahlfors regular measures are always uniformly perfect. For such measures, it was shown by Bortz et. al. [10, Lemma 2.1] that it is possible to "localise" the measure without losing Ahlfors regularity. We do not know if a similar localisation is possible for uniformly perfect measures.
@@ 2.4. Further auxiliary lemmas
To prove the first claim, note that $\mathrm{Bor}(\mathbb{R}^{2d})=\mathrm{Bor}(\mathbb{R}^{d})\times\mathrm{Bor}(\mathbb{R}^{d})$ is the $\sigma$-algebra generated by the $\pi$-system of rectangles $A\times B$, $A,B\in\mathrm{Bor}(\mathbb{R}^{d})$. The two measures $f\mu\times g\sigma$ and $(f\times g)(\mu\times\sigma)$ clearly agree on this $\pi$-system, and have common (finite) mass. So, by either Dynkin’s lemma or the monontone class Lemma [14, Lemma 2.35], they agree on $\mathrm{Bor}(\mathbb{R}^{2d})$. ∎
@@ 3. Main technical proposition
We start by introducing further notation. Let $\psi=\psi_{d}\in C^{\infty}_{c}(\mathbb{R}^{d})$ be a fixed radially decreasing function satisfying $\int\psi=1$ and $\mathbf{1}_{B(1/2)}\leq\psi\lesssim\mathbf{1}_{B(1)}$. Starting from $\psi$, define the standard "approximate identity" family $\{\psi_{\delta}\}_{\delta>0}$, where $\psi_{\delta}(x)=\delta^{-d}\psi(x/\delta)$.
For $\delta\in(0,1]$, $d\in\mathbb{N}$, $\alpha\in(0,d)$, and a Radon measure $\mu$ on $\mathbb{R}^{d}$, the notation $I_{\alpha}^{\delta}(\mu)$ refers to the $\alpha$-dimensional Riesz-energy of the mollified measure $\mu_{\delta}=\mu\ast\psi_{\delta}$. Thus,
$I_{\alpha}^{\delta}(\mu):=I_{\alpha}(\mu_{\delta}):=\iint\frac{\mu_{\delta}(x)\mu_{\delta}(y)}{|x-y|^{\alpha}}\,dx\,dy.$
@@
    $\operatorname{spt}\mu$ is contained in a dyadic cube of side length $1$, $\mu(\mathbb{R}^{2})\leq 1$, and $I_{\alpha}^{\delta}(\mu)\leq\delta^{-\epsilon}$;
@@
Note that the exponent $\epsilon$ is independent of the diameter constant $\mathfrak{d}$.
@@
$\epsilon:=\epsilon(\alpha,\beta):=\tfrac{1}{30}\min\{\epsilon_{0},2-\alpha\}.$ (3.5)
@@
Given $\rho\in 2^{-\mathbb{N}}$, let $\mu_{\rho}$ be the restriction of $\mu$ to those squares $p\in\mathcal{D}_{\delta}(Q)$ such that
@@
$(\mu_{\rho}\times\sigma)(G)\geq\delta^{2\epsilon}.$ (3.6)
Indeed, this follow since $\mu=\sum_{\rho}\mu_{\rho}$, by dyadic pigeonholing, and by assuming $\delta>0$ is sufficiently small in terms of $\epsilon$.
@@
$\mathcal{P}:=\mathcal{D}_{\delta}(\operatorname{spt}\mu_{\rho}).$
@@
$\mu_{\rho}(\cup\mathcal{R})\lesssim\delta^{3\epsilon}.$
@@
$(\mu_{\rho}\times\sigma)((\cup\mathcal{R}\times\mathbb{R}^{2})\cap G)\lesssim\delta^{3\epsilon},$
@@
$(\mu_{\rho}\times\sigma)((\cup\mathcal{R}\times\mathbb{R}^{2})\cap G)\leq\tfrac{1}{2}(\mu_{\rho}\times\sigma)(G).$
@@
$(\mu_{\rho}\times\sigma)((X\times\mathbb{R}^{2})\cap G)\gtrsim\delta^{6\epsilon}(\mu_{\rho}\times\sigma)(G)\geq\delta^{8\epsilon}.$ (3.8)
Since $I_{\alpha}^{\delta}(\mu)\leq\delta^{-\epsilon}$ by hypothesis, and $\mu(X)\geq\mu_{\rho}(X)\gtrsim\delta^{8\epsilon}$ (by (3.8)), we may infer (for $\delta>0$ small enough) from [22, Lemma 4.1] that
$|\mathcal{X}|\geq\delta^{-\alpha+26\epsilon}.$ (3.9)
@@
It follows from (3.8), and from the density constancy $\mu_{\rho}(p)/\nu(p)\sim\rho/|\mathcal{X}|$ for $p\in\mathcal{X}$, that for $\nu$ the uniform probability on $X$ (as in Proposition 3.3 Part (3)),
@@
Since $\epsilon_{0}-26\epsilon\geq\epsilon$ by (3.5), we we have shown that $|E|_{\delta}\geq\delta^{-\alpha-\epsilon}$ in all cases. The proof is complete. ∎
@@ 3.1.3. Slices and projections of our chosen cube
4. Energy flattening, and proof of Theorem 1.2 and Corollary 1.3
In this section we complete the proof of Theorem 1.2. With Proposition 3.1 in hand, the argument is adapted from the deduction of [22, Theorem 1.1] from [22, Proposition 4.3].
We first use Proposition 3.1 to find an "energy flattening" statement of roughly the following kind: if $\sigma$ is a $(D,\beta)$-uniformly perfect measure on $\mathbb{P}$, and $\mu$ is a probability measure on $\mathbb{R}^{2}$ with finite $\alpha$-energy, $\alpha<2$, then $(\mu\ast\sigma)^{k}$ has finite $(\alpha+\eta)$-energy for some $\eta=\eta(\alpha,\beta,D)>0$, and provided $k\in\mathbb{N}$ is large enough. The precise statement is Proposition 4.2. Applying this result iteratively to $\mu=\sigma^{k}$ eventually shows that $\sigma^{k}$ has finite $t$-energy for index $t$ arbitrarily close to $2$. This is detailed in Corollary 4.6. From this Theorem 1.2 follows by a short argument, see Section 4.2.
4.1. Energy flattening
@@
    $\mu$ is supported a dyadic cube of side length $1$, $\mu(\mathbb{R}^{2})\leq 1$, and $I_{\alpha}^{\delta}(\mu)\leq\delta^{-\epsilon}$;
@@
Next, we prove the uniformity in $\epsilon$ as in the last assertion of the Corollary. As we have just seen, there is no loss of generality in assuming $k=1$, so we focus on this case. Fix $\beta,D$ as in Proposition 3.1, and let $I\subset(0,2)$ be a compact interval. We have already proved Proposition 3.1 is valid for each $\alpha\in I$. Let $\epsilon_{\alpha}:=\epsilon(\alpha,\beta,D)>0$ be the constant produced by the proposition. The open intervals $B(\alpha,\epsilon_{\alpha}/4),\alpha\in I,$ cover $I$, so by compactness we may choose a finite subset $\mathcal{A}\subset I$ such that the intervals $\{B(\alpha,\epsilon_{\alpha}/4):\alpha\in\mathcal{A}\}$ already cover $I$. Set
@@
Let us check that $\mu,\sigma,E$ satisfy hypotheses (1)-(3) with constants $(\alpha^{\prime},\epsilon_{\alpha^{\prime}})$. Regarding (1), it follows from the Fourier-analytic expression $I_{s}(\nu)=c(s)\int|\hat{\nu}(\xi)|^{2}|\xi|^{s-d}\,d\xi$ for the $s$-energy [18, Lemma 12.12], and then from the definitions above, that
$I_{\alpha^{\prime}}^{\delta}(\mu)\lesssim_{I}\delta^{-2|\alpha-\alpha^{\prime}|}I_{\alpha}(\mu)+\delta\leq\delta^{-\epsilon_{\alpha^{\prime}}/2-\epsilon}+\delta\leq\delta^{-3\epsilon_{\alpha^{\prime}}/4}+\delta.$
The displayed inequality above shows that, for $\delta>0$ small enough depending only on $I,\mathcal{A}$, it holds that
$I_{\alpha^{\prime}}^{\delta}(\mu)\leq\delta^{-\epsilon_{\alpha^{\prime}}}.$
This is what we need for (1). Part (2) holds trivially, and part (3) follows from our choice of $\epsilon$ (4.1), as
@@
Proposition 4.2 below is based on [22, Proposition 4.7], and the proof is virtually the same, except that we rely on Corollary 3.1 instead of [22, Proposition 4.3]. Our proof also eschews a parabolic rescaling argument [22, Proposition 4.7] (it turns out that one can get rid of this argument by a little additional pigeonholing).
@@
For all $\alpha\in[0,2)$, $\beta\in[0,1)$, $D>1$, $\mathfrak{d}>0$, $\kappa>0$, and $R>1$ there exist constants
$\begin{cases}\epsilon=\epsilon(\alpha,\beta,\kappa,D)>0,\\ k_{0}=k_{0}(\alpha,\beta,\kappa,D)\in\mathbb{N},\\ \delta_{0}=\delta_{0}(\alpha,\beta,\kappa,\mathfrak{d},D,R)>0,\\ \eta=\eta(\alpha,\beta,D)>0\end{cases}$
@@
    $\mu(\mathbb{R}^{2})\leq 1$, $\operatorname{spt}\mu\subset[-R,R)^{2}$, and $I_{\alpha}^{\delta}(\mu)\leq\delta^{-\epsilon}$;
@@
$I_{\alpha+\eta}^{\delta}\left(\left(\mu\ast\sigma\right)^{k}\right)\leq\delta^{-\kappa},\qquad k\geq k_{0}.$
@@
In the following the implicit constants in the $\lesssim"$ may depend on the parameters $D,\alpha,\beta,\kappa$. We start by the parameters $\eta,\epsilon,k_{0}$ whose existence is claimed. We first set
$\eta:=\min\{2-\alpha,\epsilon(\alpha,\beta,D)\},$
where $\epsilon(\alpha,\beta,D)>0$ is the constant output by Corollary 4.1. Next, we set
$\kappa^{\prime}:=\min\{\eta,\kappa\}\quad\text{and}\quad\epsilon:=\frac{\kappa^{\prime}\eta}{10}\leq\frac{\kappa\eta}{10}\quad\text{and}\quad k_{0}:=2^{\lceil\tfrac{20}{\kappa^{\prime}}\rceil}.$ (4.2)
We then set up some further notation. For $r>0$ and $k\in\mathbb{N}$, we denote
$\Pi_{r}^{k}:=\left(\mu\ast\sigma\right)^{k}\ast\psi_{r},\text{ and }J_{r}(k):=\left\lVert\Pi_{r}^{2^{k}}\right\rVert_{2}.$
We record the following simple reduction.
@@
Let $k_{0}=\lceil 20/\kappa^{\prime}\rceil$, and suppose that
$J_{r}(k_{0})\leq\delta^{-\frac{\kappa}{4}}r^{\frac{\alpha+\eta-2}{2}},\quad\delta\leq r\leq 1.$ (4.3)
Then the conclusion of Proposition 4.2 holds.

### v2

@@ Abstract.
Fourier decay, inverse theorems, sumsets, convolutions.
@@ 1. Introduction
Theorem 1.2 can be viewed as an $L^{2}$-flattening statement: upon mollifying the measure, iterated self-convolutions yield quantitative decay of its $L^{2}$ norm; see, for instance, Proposition 4.2 or Corollary 4.5 below. For further discussion of the relation between (1.3) and flattening of (various notions of) $L^{2}$ norms, see [2, Corollary 1.2].
@@ 1.1. Previous work
Close relatives of Theorem 1.2 in previous literature are the $L^{2}$-flattening theorems of Rossi-Shmerkin [25, Theorem 1.1] on $\mathbb{R}$, and Khalil [18, Theorem 1.6] on $\mathbb{R}^{d}$. These results are formulated in terminology different from Theorem 1.2, but they are roughly the counterparts of Theorem 1.2 for (i) uniformly perfect measures on $\mathbb{R}$, and (ii) measures on $\mathbb{R}^{d}$ satisfying Khalil’s uniformly affine non-concentration condition. The reader should note that measures supported on smooth curves – as in Theorem 1.2 – do not satisfy Khalil’s non-concentration condition, due to the presence of tangent lines; otherwise Theorem 1.2 could be deduced from the results in [18].
Thus, Theorem 1.2 and [18, Theorem 1.6] are complementary, but finding a (natural) common generalisation seems like an interesting problem. For expert readers, we also mention that existing higher-dimensional inverse theorems (by Hochman [17] and Shmerkin [27]) do not appear to be powerful enough to prove Theorem 1.2, again due to the existence of tangent lines. Our proof will eventually rely on the one-dimensional inverse theorem of Shmerkin [26, Theorem 2.1], see Section 1.2 for a brief explanation.
@@
Finally, Theorem 1.2 is related to several recent works studying Fourier decay of stationary measures. The first author and Khalil [2] proved that the conclusion of Theorem 1.2 holds under the following assumptions. The measure $\sigma$ is the lift of a non-atomic self-similar measure on $\mathbb{R}$, onto either (a) an analytic curve whose trace is not contained in an affine hyperplane of $\mathbb{R}^{d}$, or (b) a $C^{d+1}$-curve $\gamma$ such that $\{\gamma^{\prime},\gamma^{\prime\prime},...,\gamma^{(d)}\}$ span $\mathbb{R}^{d}$ at every point. This was the first demonstration of a general non-trivial class of measures on curves for which $L^{2}$ flattening can be obtained, in the sense (1.3). This work is also related to that of Algom, Chang, Meng Wu, and Yu-Liang Wu [1], and the simultaneous independent work by Baker and Banaji [7], on pointwise Fourier decay for smooth strictly convex push-forwards of self-similar measures. It is also related to the subsequent paper of Baker, Khalil, and Sahlsten [8], as well as to [3, 4, 5, 9, 10, 12, 19, 28, 29, 30]. For more on the relation between Theorem 1.2 and the Fourier decay problem for stationary measures, see [2, Section 1.2].
@@
It is also natural to consider analogues of Theorem 1.2 for surfaces. For instance, one may ask whether a Borel probability measure $\sigma$ on the paraboloid (or any other "curved" hypergraph) in $\mathbb{R}^{d+1}$ is $L^{2}$ flattening, provided its projection to $\mathbb{R}^{d}$ is uniformly affinely non-concentrated in the sense of Khalil [18]. More generally, it may be plausible that $\sigma$ is $L^{2}$ flattening whenever its projection to $\mathbb{R}^{d}$ is $L^{2}$ flattening. A strategy like this underlies the proof of $L^{2}$ flattening for self-similar measures on curves, cf. [2, Theorem 1.3].
@@ 1.2. Proof outline
Our proof is based on Shmerkin’s inverse theorem for $L^{q}$-norms [26, Theorem 2.1], which in turn is inspired by Hochman’s inverse theorem for entropy [16, Theorem 2.7]. In fact, we employ a corollary of Shmerkin’s theorem obtained by Rossi and Shmerkin [25, Proposition 3.1], which states that convolution with a uniformly perfect measure on the real line is $L^{2}$-flattening. One may think of $Y$ in the "Lemma" as the support of our uniformly perfect measure $\sigma$ in Theorem 1.2, although this is slightly misleading; Lemma 3.2 deals directly with the measure $\sigma$, not its support. The support analogy is more accurate if $\sigma$ happens to be Ahlfors regular (and Figure 1 depicts this case).
@@
From Proposition 3.3 (and a "measure-theoretic" version of it stated in Proposition 3.1) adapting a strategy developed in [21, 22], we derive Proposition 4.2: convolution with a uniformly perfect measure $\sigma$ on $\mathbb{P}$ is $L^{2}$-flattening. In particular, taking repeated self-convolutions of $\sigma$ gradually and quantitatively decreases the $L^{2}$-norm of the convolution. This is formalised in Corollary 4.5. This easily yields Theorem 1.2, see Section 4.2.
@@ 2.2. Rossi and Shmerkin’s theorem
We need this refined definition since in our application of Rossi and Shmerkin’s theorem, we shall require certain restrictions of the original measure to be uniformly perfect (see Lemma 2.14 below). Now, by [25, Lemma 4.1], Ahlfors regular measures are always uniformly perfect. For such measures, it was shown by Bortz et. al. [11, Lemma 2.1] that it is possible to "localise" the measure without losing Ahlfors regularity. We do not know if a similar localisation is possible for uniformly perfect measures.
@@ 2.4. Further auxiliary lemmas
To prove the first claim, note that $\mathrm{Bor}(\mathbb{R}^{2d})=\mathrm{Bor}(\mathbb{R}^{d})\times\mathrm{Bor}(\mathbb{R}^{d})$ is the $\sigma$-algebra generated by the $\pi$-system of rectangles $A\times B$, $A,B\in\mathrm{Bor}(\mathbb{R}^{d})$. The two measures $f\mu\times g\sigma$ and $(f\times g)(\mu\times\sigma)$ clearly agree on this $\pi$-system, and have common (finite) mass. So, by either Dynkin’s lemma or the monontone class Lemma [15, Lemma 2.35], they agree on $\mathrm{Bor}(\mathbb{R}^{2d})$. ∎
@@ 3. Main technical proposition
We start by introducing further notation. Let $\psi=\psi_{d}\in C^{\infty}_{c}(\mathbb{R}^{d})$ be a fixed radially decreasing function satisfying $\int\psi=1$ and $\mathbf{1}_{B(1/2)}\leq\psi\lesssim\mathbf{1}_{B(1)}$. Starting from $\psi$, define the standard "approximate identity" family $\{\psi_{\delta}\}_{\delta>0}$, where $\psi_{\delta}(x)=\delta^{-d}\psi(x/\delta)$. For a Radon measure $\mu$ on $\mathbb{R}^{d}$, we write $\mu_{\delta}:=\mu\ast\psi_{\delta}$.
@@
    $\operatorname{spt}\mu$ is contained in a dyadic cube of side length $1$, $\mu(\mathbb{R}^{2})\leq 1$, and $\|\mu_{\delta}\|_{2}^{2}\leq\delta^{\alpha-2-\epsilon}$;
@@
Note that the exponent $\epsilon$ is independent of the diameter constant $\mathfrak{d}$. The notation $\|\cdot\|_{2}$ refers to the usual $L^{2}$-norm (as opposed to the notation $\|\cdot\|_{L^{2},\mathrm{Sh}}$ from Definition 2.3).
@@
$\epsilon:=\epsilon(\alpha,\beta):=\tfrac{1}{20}\min\{\epsilon_{0},2-\alpha\}.$ (3.5)
@@
Given $\rho\in 2^{-\mathbb{N}}$, let $\mu^{\rho}$ be the restriction of $\mu$ to those squares $p\in\mathcal{D}_{\delta}(Q)$ such that
@@
$(\mu^{\rho}\times\sigma)(G)\geq\delta^{2\epsilon}.$ (3.6)
Indeed, this follow since $\mu=\sum_{\rho}\mu^{\rho}$, by dyadic pigeonholing, and by assuming $\delta>0$ is sufficiently small in terms of $\epsilon$.
@@
$\mathcal{P}:=\mathcal{D}_{\delta}(\operatorname{spt}\mu^{\rho}).$
@@
$\mu^{\rho}(\cup\mathcal{R})\lesssim\delta^{3\epsilon}.$
@@
$(\mu^{\rho}\times\sigma)((\cup\mathcal{R}\times\mathbb{R}^{2})\cap G)\lesssim\delta^{3\epsilon},$
@@
$(\mu^{\rho}\times\sigma)((\cup\mathcal{R}\times\mathbb{R}^{2})\cap G)\leq\tfrac{1}{2}(\mu^{\rho}\times\sigma)(G).$
@@
$(\mu^{\rho}\times\sigma)((X\times\mathbb{R}^{2})\cap G)\gtrsim\delta^{6\epsilon}(\mu^{\rho}\times\sigma)(G)\geq\delta^{8\epsilon}.$ (3.8)
Since $\|\mu_{\delta}\|_{2}^{2}\leq\delta^{\alpha-2-\epsilon}$ by hypothesis, and $\mu_{\delta}(X)\geq\mu^{\rho}(X)\gtrsim\delta^{8\epsilon}$ (by (3.8)), we may infer from Cauchy-Schwarz that
$\delta^{8\epsilon}\lesssim\int_{X}\mu_{\delta}\leq(\mathrm{Leb}(X))^{1/2}\delta^{(\alpha-2-\epsilon)/2}\quad\Longrightarrow\quad\mathrm{Leb}(X)\gtrsim\delta^{2-\alpha+17\epsilon}.$
Consequently, provided $\delta>0$ is small enough in terms of $\epsilon$,
$|\mathcal{X}|\geq\delta^{-\alpha+18\epsilon}.$ (3.9)
@@
Let $\nu$ be the uniform probability on $X$ (as in Proposition 3.3 Part (3)). It follows from (3.8), and from the density constancy $\mu_{\rho}(p)/\nu(p)\sim\rho/|\mathcal{X}|$ for $p\in\mathcal{X}$, that
@@
Since $\epsilon_{0}-18\epsilon\geq\epsilon$ by (3.5), we we have shown that $|E|_{\delta}\geq\delta^{-\alpha-\epsilon}$ in all cases. The proof is complete. ∎
@@ 3.1.3. Slices and projections of our chosen cube
4. $L^{2}$-flattening and proof of Theorem 1.2
In this section we complete the proof of Theorem 1.2. With Proposition 3.1 in hand, the argument is adapted from the deduction of [22, Theorem 1.1] from [22, Proposition 4.3], except that we work with $L^{2}$ norms rather than energies.
We first apply Proposition 3.1 to obtain an $L^{2}$-flattening statement of roughly the following form: if $\sigma$ is a $(D,\beta)$-uniformly perfect measure on $\mathbb{P}$ and $\mu$ is a probability measure on $\mathbb{R}^{2}$, then the $L^{2}$ norm of $(\mu\ast\sigma)^{k}_{\delta}$ is smaller than that of $\mu_{\delta}$ by a factor of $\delta^{\eta}$, provided $k\in\mathbb{N}$ is sufficiently large. The precise formulation is Proposition 4.2. Applying this result iteratively shows that the $L^{2}$ norm of $\sigma_{\delta}^{k}$ can be made $\leq\delta^{-\epsilon}$ by choosing $k=k(\epsilon)\geq 1$ sufficiently large. The details are carried out in Corollary 4.5. From there, Theorem 1.2 follows easily, see Section 4.2.
4.1. $L^{2}$-flattening
@@
    $\mu$ is supported on a dyadic cube of side length $1$, $\mu(\mathbb{R}^{2})\leq 1$, and $\|\mu_{\delta}\|_{2}^{2}\leq\delta^{\alpha-2-\epsilon}$;
@@
Next, we prove the uniformity in $\epsilon$ as in the last assertion of the Corollary. As we have just seen, there is no loss of generality in assuming $k=1$, so we focus on this case. Fix $\beta,D$ as in Proposition 3.1, and let $I\subset(0,2)$ be a compact interval, and fix $\bar{\alpha}\in(\max I,2)$. We have already proved that Proposition 3.1 is valid for each $\alpha\in I$. Let $\epsilon_{\alpha}:=\epsilon(\alpha,\beta,D)>0$ be the constant produced by the proposition. The open intervals $B(\alpha,\epsilon_{\alpha}/4),\alpha\in I,$ cover $I$, so by compactness we may choose a finite subset $\mathcal{A}\subset I$ such that the intervals $\{B(\alpha,\epsilon_{\alpha}/4):\alpha\in\mathcal{A}\}$ already cover $I$. Set
@@
Let us check that $\mu,\sigma,E$ satisfy hypotheses (1)-(3) with constants $(\alpha^{\prime},\epsilon_{\alpha^{\prime}})$. Regarding (1),
$\|\mu_{\delta}\|_{2}^{2}\leq\delta^{\alpha-2-\epsilon}\leq\delta^{\alpha^{\prime}-2-\epsilon-|\alpha-\alpha^{\prime}|}\leq\delta^{\alpha^{\prime}-2-\epsilon_{\alpha^{\prime}}}.$
Part (2) holds trivially, and part (3) follows from our choice of $\epsilon$ (4.1), as
@@
Proposition 4.2 below is modeled on [22, Proposition 4.7]. The proof, however, differs in that we work with $L^{2}$ norms instead of energies, and in particular we invoke Corollary 4.1 in place of [22, Proposition 4.3]. Our argument also avoids the parabolic rescaling step in [22, Proposition 4.7]; it turns out that this can be dispensed with at the cost of a little additional pigeonholing.
@@
For all $\alpha\in[0,2)$, $\beta\in[0,1)$, $D>1$, $\mathfrak{d}>0$, and $R>1$ there exist constants
$\begin{cases}\eta=\eta(\alpha,\beta,D)>0,\\ k_{0}=k_{0}(\alpha,\beta,D)\in\mathbb{N},\\ \delta_{0}=\delta_{0}(\alpha,\beta,\mathfrak{d},D,R)>0,\end{cases}$
@@
    $\mu(\mathbb{R}^{2})\leq 1$, $\operatorname{spt}\mu\subset[-R,R)^{2}$, and $\|\mu_{\delta}\|_{2}^{2}\leq\delta^{\alpha-2}$;
@@
$\big\|(\mu\ast\sigma)_{\delta}^{k}\big\|_{2}^{2}\leq\delta^{\alpha+\eta-2},\qquad k\geq k_{0}.$
@@
In the following the implicit constants in the $\lesssim"$ may depend on the parameters $D,\alpha,\beta,\kappa$. We start by defining the parameters $\eta,\epsilon,k_{0}$ whose existence is claimed. We set
$\eta:=\min\{\tfrac{1}{2}\epsilon(\alpha,\beta,D),2-\alpha\}\quad\text{and}\quad k_{0}:=2^{\lceil 20/\eta\rceil},$ (4.2)
where $\epsilon(\alpha,\beta,D)>0$ is the constant given by Corollary 4.1. We then set up some further notation. For $k\in\mathbb{N}$, we denote
$\Pi^{k}:=\Pi^{k}_{\delta}:=\left(\mu\ast\sigma\right)^{k}\ast\psi_{\delta},\text{ and }J(k):=\big\|\Pi^{2^{k}}\big\|_{2}.$
We record that the sequence $J(k)$ is non-increasing by Young’s inequality (or Plancherel), writing $\left\lVert\nu\right\rVert_{1}$ for both total variance and $L^{1}$-norm:
$J(k+1)=\big\|\Pi^{2^{k}}\ast\Pi^{2^{k}}\big\|_{2}\leq\big\|\Pi^{2^{k}}\big\|_{1}\cdot\big\|\Pi^{2^{k}}\big\|_{2}=\big\|\Pi^{2^{k}}\big\|_{2}=J(k),\quad k\geq 0.$
In particular, to prove Proposition 4.2, it suffices to find $k\leq\lceil 20/\eta\rceil$ such that
$J(k)\leq\delta^{\frac{\alpha+\eta-2}{2}}.$ (4.3)
@@
There either exists $k\leq\lceil 20/\eta\rceil$ such that
$\delta^{\eta/10}J(k)\leq J(k+1)\leq J(k),$ (4.4)
or otherwise (4.3) holds with $k=\lceil 20/\eta\rceil$ (provided $\delta>0$ is sufficiently small in terms of $R$).